OpenStudy (anonymous):

Julio is drawing tiles from a bag of tiles like the one shown below. He draws a tile, replaces it, and draws another tile. What is the probability he will draw a red tile on both draws? 1/25 16/225 2/21 1/9

OpenStudy (anonymous):

@Mason_Jones573

OpenStudy (anonymous):

@ganeshie8 @Loser66

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

There are five red and a total of fifteen tiles. So the probability of getting a red tile on the first attempt is 5/15 = 1/3 Since Julio replaces it before drawing the next time, so the probabilities do not change, and hence getting a red tile is 1/3. The two events - drawing a tile each time - are independent of each other and hence the probability of getting two tiles is a product of the two individual probabilities, which is\[ 1/3 * 1/3 = 1/9 \]